# Longitudinal efficient adjustment sets for time-varying treatment effect estimation in nonparametric causal graphical models

### v1

@@ 4 Comparison of time-dependent adjustment sets
$\mathbf{G}_{k}\perp\!\!\!\perp\mathbf{B}_{k-1}\ \mid\ \mathbf{G}_{k-1},\mathbf{A}_{k-1}\ for\ k=1,...,p.$ (5)
Then $\mathbf{G}=(\mathbf{G}_{0},\mathbf{G}_{1},...,\mathbf{G}_{p})$ is also a sufficient time-dependent adjustment set according to Definition 2 and
$\sigma^{2}_{\mathbf{a},\mathbf{G},\mathbf{B}}(P)-\sigma^{2}_{\mathbf{a},\mathbf{G}}(P)\geq 0\ \ and\ \ \sigma^{2}_{\Delta,\mathbf{G},\mathbf{B}}(P)-\sigma^{2}_{\Delta,\mathbf{G}}(P)\geq 0.$
@@
Let $\mathbf{K}=(\mathbf{K}_{0},\mathbf{K}_{1},...,\mathbf{K}_{p})$ be a sufficient time-dependent adjustment set according to Definition 1 and let ${\mathbf{B}_{k}\subseteq\{\bar{\mathbf{A}}_{k-1},\bar{\mathbf{K}}_{k}\}}$ and ${\mathbf{G}_{k}\subseteq\{\bar{\mathbf{A}}_{k-1},\bar{\mathbf{K}}_{j}\}\setminus\mathbf{B}_{k}},k=0,...,p$ be two disjoint subsets of $\{\bar{\mathbf{A}}_{j-1},\bar{\mathbf{K}}_{k}\}$ such that $(\mathbf{G}_{k},\mathbf{B}_{k})=\{\bar{\mathbf{A}}_{k-1},\bar{\mathbf{K}}_{k}\}$. Suppose
$\mathbf{B}=(\mathbf{B}_{0},\mathbf{B}_{1},...,\mathbf{B}_{p})$
@@
$\mathbf{G}=(\mathbf{G}_{0},\mathbf{G}_{1},...,\mathbf{G}_{p})$
@@
are two sufficient time dependent adjustment sets according to Definition 2. Suppose that
$A_{k}\perp\!\!\!\perp(\mathbf{G}_{k}\setminus\mathbf{B}_{k})\ \mid\ \mathbf{B}_{k},\ for\ k=0,...,p,$
$Y\perp\!\!\!\perp(\mathbf{B}_{p}\setminus\mathbf{G}_{p})\ \mid\ \mathbf{G}_{p},\mathbf{A}_{p}.$
and
$\mathbf{G}_{k}\perp\!\!\!\perp(\mathbf{B}_{k-1}\setminus\mathbf{G}_{k-1})\ \mid\ \mathbf{G}_{k-1},\mathbf{A}_{k-1}\ for\ k=1,...,p.$
Then for all distributions $P$ compatible with the causal graph $\mathcal{G}$
$\sigma^{2}_{\mathbf{a},\mathbf{B}}(P)-\sigma^{2}_{\mathbf{a},\mathbf{G}}(P)\geq 0\ \ and\ \ \sigma^{2}_{\Delta,\mathbf{B}}(P)-\sigma^{2}_{\Delta,\mathbf{G}}(P)\geq 0.$
@@
We now return to Example 1 once more to illustrate how these lemmas and theorem can be used in practice to identify a time-dependent adjustment set yielding an estimator with reduced variance. Next, we revisit Example 4 of Rotnitzky and Smucler (2020) [5]. In this example, an optimal adjustment set cannot be identified based on the causal graph alone when considering Definition 1 and associated lemmas and theorem. However, our numerical example suggests that our extended results may resolve this problem.
@@
Table 1 lists all time-dependent adjustment sets according to Definition 2. Note that only adjustment sets 1-9 satisfy Definition 1. Based on Theorem 4.3, the asymptotic variance of estimators based on sets 5 and 14 can be compared, for example. Let $\mathbf{B}=(\mathbf{B_{0}}=C_{0,2},\mathbf{B_{1}}=\{C_{1,2},A_{0},C_{0,2}\})$ and $\mathbf{G}=(\mathbf{G_{0}}=C_{0,2},\mathbf{G_{1}}=\{C_{1,2},A_{0}\})$.Because $A_{1}\perp\!\!\!\perp\varnothing\mid\mathbf{B_{1}}$, $A_{0}\perp\!\!\!\perp\varnothing\mid\mathbf{B_{0}}$, $Y\perp\!\!\!\perp C_{0,2}\mid A_{1},C_{1,2},A_{0}$, and $(C_{1,2},C_{0})\perp\!\!\!\perp\varnothing\mid C_{0,2}A_{0}$, we know that the asymptotic variance of nonparametric efficient estimators based on set 14 is lower or equal to the asymptotic variance of nonparametric efficient estimators based on set 5. We provide a numerical illustration based on simulated data to support these theoretical results. We simulated 10,000 datasets of size $n=$ 1,000 using the following data generating equations:
$C_{01}$ $\sim TN(\mu=0,\sigma^{2}=1,\min=-2,\max=2)$
$C_{02}$ $\sim C_{01}+TN(\mu=0,\sigma^{2}=1,\min=-2,\max=2)$
$A_{0}$ $\sim Bernouilli(p=\text{expit}(C_{01}))$
$C_{11}$ $\sim A_{0}+C_{02}+TN(\mu=0,\sigma^{2}=1,\min=-2,\max=2)$
$C_{12}$ $\sim C_{11}+TN(\mu=0,\sigma^{2}=1,\min=-2,\max=2)$
$Y$ $\sim A_{1}+A_{0}+C_{12}+TN(\mu=0,\sigma^{2}=1,\min=-2,\max=2),$
where $TN$ refers to a truncated normal distribution (in order to avoid practical positivity violations) and $\text{expit}(\cdot)=\exp(\cdot)/(1+\exp(\cdot))$ is the inverse of the logit function. The number of replications was chosen to ensure that the Monte Carlo standard error for the standard deviation was $\approx$ 0.001, which is sufficiently small to accurately identify the adjustment set with the lowest variance. We estimated $\mathbb{E}(Y^{11})$ using a pooled longitudinal targeted maximum likelihood estimator, which is based on the nonparametric efficient influence function $\psi_{P,\mathbf{a}}(\mathbf{Z};\mathcal{G})$. This was achieved in R v4.3.2 using the ltmle package v2.0.0 and generalized linear models to model both the outcome and exposure.
@@
Adjustment set $\mathbf{Z_{0}}$ $\mathbf{Z_{1}}$ Monte Carlo
@@
As a second example, we revisit Example 4 from Rotnitzky and Smucler (2020) [5]. The associated causal graph is represented in Figure 3. In this example, the optimal sufficient adjustment set cannot be uniquely identified from the causal graph alone when considering Definition 1. Indeed, as illustrated in Rotnitzky and Smucler (2020) [5], the optimal sufficient adjustment set is either 1 or 8 depending on the data-generating equations, more specifically on the relative strength of the $H\rightarrow R\rightarrow Q$ and $H\rightarrow A_{1}$ pathways.
@@
Table 2 lists all sufficient adjustment sets according to Definition 2. Only sets 1-11 satisfy Definition 1. Using Theorem 4.3, it can be found that estimators based on set 24 have a lower variance than those based on either sets 1 and 8. For example, Theorem 4.3 allows comparing sets 1 and 24 by letting $\mathbf{B}=(\mathbf{B_{0}}=\varnothing,\mathbf{B_{1}}=\{A_{0},Q\}),\mathbf{G}=(\mathbf{G_{0}}=H,\mathbf{G_{1}}=Q)$ and noticing that $A_{1}\perp\!\!\!\perp\varnothing\mid\mathbf{B_{1}}$, $A_{0}\perp\!\!\!\perp H\mid\varnothing$, $Y\perp\!\!\!\perp A_{0}\mid Q,A_{1}$ and $Q\perp\!\!\!\perp\varnothing\mid H,A_{0}$. Similarly, sets 8 and 24 can be compared by letting $\mathbf{B}=(\mathbf{B_{0}}=H,\mathbf{B_{1}}=\{A_{0},Q,H\}),\mathbf{G}=(\mathbf{G_{0}}=H,\mathbf{G_{1}}=Q)$ and noticing that $A_{1}\perp\!\!\!\perp\varnothing\mid\mathbf{B_{1}}$, $A_{0}\perp\!\!\!\perp\varnothing\mid\mathbf{B_{0}}$, $Y\perp\!\!\!\perp(A_{0},H)\mid Q,A_{1}$ and $Q\perp\!\!\!\perp\varnothing\mid H,A_{0}$. To illustrate empirically these results, we generated 20,000 datasets of size $n$ = 1,000 according to the following equations
@@
Adjustment set $\mathbf{Z_{0}}$ $\mathbf{Z_{1}}$ Monte Carlo SD
@@
1 $\varnothing$ $\{A_{0},Q\}$ 0.146 0.199
2 $\varnothing$ $\{A_{0},R\}$ 0.175 0.222
3 $\varnothing$ $\{A_{0},H\}$ 0.214 0.248
4 $\varnothing$ $\{A_{0},Q,R\}$ 0.154 0.203
5 $\varnothing$ $\{A_{0},Q,H\}$ 0.165 0.207
6 $\varnothing$ $\{A_{0},R,H\}$ 0.191 0.229
7 $\varnothing$ $\{A_{0},Q,R,H\}$ 0.165 0.207
@@
12 $\varnothing$ $Q$ 0.139 0.194
13 $\varnothing$ $R$ 0.158 0.213
14 $\varnothing$ $\{Q,R\}$ 0.143 0.197
15 $\varnothing$ $\{Q,H\}$ 0.165 0.208
16 $\varnothing$ $\{R,H\}$ 0.191 0.229
17 $\varnothing$ $\{Q,R,H\}$ 0.165 0.208

### v2

@@ 4 Comparison of sufficient time-dependent adjustment sets
$Q_{t-1}\perp\!\!\!\perp_{\mathcal{G}_{t-1}}\boldsymbol{B}_{t-1},\boldsymbol{D}_{t-1}\ \mid\ \boldsymbol{G}_{t-1},\boldsymbol{C}_{t-1},{A}_{t-1}\ for\ t=T,...,1.$ (6)
Then, for any distribution $P$ compatible with the graph $\mathcal{G}$, $(\boldsymbol{G},\boldsymbol{C})=\{(\boldsymbol{G}_{0},\boldsymbol{C}_{0}),(\boldsymbol{G}_{1},\boldsymbol{C}_{1}),\dots,(\boldsymbol{G}_{T},\boldsymbol{C}_{T})\}$ is also a sufficient time-dependent adjustment set according to Definition 2 and
  1. 1.
    $\sigma^{2}_{\boldsymbol{a},\boldsymbol{G},\boldsymbol{B},\boldsymbol{C},\boldsymbol{D}}(P)-\sigma^{2}_{\boldsymbol{a},\boldsymbol{G},\boldsymbol{C}}(P)\geq 0\ \ and\ \ \sigma^{2}_{\Delta,\boldsymbol{G},\boldsymbol{B},\boldsymbol{C},\boldsymbol{D}}(P)-\sigma^{2}_{\Delta,\boldsymbol{G},\boldsymbol{C}}(P)\geq 0.$
  2. 2.
    For any $\boldsymbol{B}^{\prime}\subseteq\boldsymbol{B}$ and $\boldsymbol{D}^{\prime}\subseteq\boldsymbol{D}$, ($\boldsymbol{G},\boldsymbol{B}^{\prime},\boldsymbol{C},\boldsymbol{D}^{\prime}$) is also sufficient time-dependent adjustment set according to Definition 2 and
    $\sigma^{2}_{\boldsymbol{a},\boldsymbol{G},\boldsymbol{B},\boldsymbol{C},\boldsymbol{D}}(P)\;\;\geq\;\;\sigma^{2}_{\boldsymbol{a},\boldsymbol{G},\boldsymbol{B}^{\prime},\boldsymbol{C},\boldsymbol{D}^{\prime}}(P)\;\;\geq\;\;\sigma^{2}_{\boldsymbol{a},\boldsymbol{G},\boldsymbol{C}}(P).$
As in Lemma 12 of Rotnitzky and Smucler [5], Lemma 4.2 states that retaining so-called superfluous adjustment variables that are associated with treatment but provide no additional predictive value for the outcome after conditioning on the existing adjustment set leads to an increase, or at least no reduction, in the asymptotic variance of the estimators.
@@
Consider a DAG $\mathcal{G}$ with vertex set $\boldsymbol{V}$. Let $\boldsymbol{A}\subset\boldsymbol{V}$ and $Y\in\boldsymbol{V}\setminus\boldsymbol{A}$. For $t=T,\dots,0$, define $\mathcal{G}_{t}$ as the subgraph of $\mathcal{G}$ constructed by applying Algorithm 1. Let $\boldsymbol{K}=(\boldsymbol{K}_{0},\boldsymbol{K}_{1},...,\boldsymbol{K}_{T})$ be a sufficient time-dependent adjustment set according to Definition 1 and let $\boldsymbol{B}_{t},\boldsymbol{G}_{t},\boldsymbol{C}_{t},\boldsymbol{D}_{t}\subseteq\{\bar{\boldsymbol{A}}_{t-1},\bar{\boldsymbol{K}}_{t}\}$, t=0,,Tt=0,...,T, be a partition of $\{\bar{\boldsymbol{A}}_{t-1},\bar{\boldsymbol{K}}_{t}\}$. Suppose $(\boldsymbol{B},\boldsymbol{C})=\{(\boldsymbol{B}_{0},\boldsymbol{C}_{0}),(\boldsymbol{B}_{1},\boldsymbol{C}_{1}),\dots,(\boldsymbol{B}_{T},\boldsymbol{C}_{T})\}$ and $(\boldsymbol{G},\boldsymbol{C})=\{(\boldsymbol{G}_{0},\boldsymbol{C}_{0}),(\boldsymbol{G}_{1},\boldsymbol{C}_{1}),\dots,(\boldsymbol{G}_{T},\boldsymbol{C}_{T})\}$ are two sufficient time-dependent adjustment sets according to Definition 2 such that
$A_{T}\perp\!\!\!\perp_{\mathcal{G}_{T}}\boldsymbol{G}_{T},\boldsymbol{D}_{T}\ \mid\ \boldsymbol{B}_{T},\boldsymbol{C}_{T},$ (7)
$Y\perp\!\!\!\perp_{\mathcal{G}_{T}}\boldsymbol{B}_{T},\boldsymbol{D}_{T}\ \mid\ \boldsymbol{G}_{T},\boldsymbol{C}_{T},{A}_{T},$ (8)
@@
$A_{t-1}\perp\!\!\!\perp_{\mathcal{G}_{t-1}}\boldsymbol{G}_{t-1},\boldsymbol{D}_{t-1}\ \mid\ \boldsymbol{B}_{t-1},\boldsymbol{C}_{t-1},$ (9)
$Q_{t-1}\perp\!\!\!\perp_{\mathcal{G}_{t-1}}\boldsymbol{B}_{t-1},\boldsymbol{D}_{t-1}\ \mid\ \boldsymbol{G}_{t-1},\boldsymbol{C}_{t-1},{A}_{t-1},\ for\ t=T,\dots,1.$ (10)
Then for all distributions $P$ compatible with the causal graph $\mathcal{G}$,
  1. 1.
    $\sigma^{2}_{\boldsymbol{a},\boldsymbol{B},\boldsymbol{C}}(P)\geq\sigma^{2}_{\boldsymbol{a},\boldsymbol{G},\boldsymbol{B},\boldsymbol{C},\boldsymbol{D}}(P)\geq\sigma^{2}_{\boldsymbol{a},\boldsymbol{G},\boldsymbol{C}}\ \ and\ \ \sigma^{2}_{\Delta,\boldsymbol{B},\boldsymbol{C}}(P)\geq\sigma^{2}_{\Delta,\boldsymbol{G},\boldsymbol{B},\boldsymbol{C},\boldsymbol{D}}(P)\geq\sigma^{2}_{\Delta,\boldsymbol{G},\boldsymbol{C}}(P).$
  2. 2.
@@
    For any $\boldsymbol{G}^{\prime}\subseteq\boldsymbol{G}$, $\boldsymbol{B}^{\prime}\subseteq\boldsymbol{B}$ and $\boldsymbol{D}^{\prime}\subseteq\boldsymbol{D}$, ($\boldsymbol{G}^{\prime},\boldsymbol{B},\boldsymbol{C},\boldsymbol{D}^{\prime}$) and ($\boldsymbol{G},\boldsymbol{B}^{\prime},\boldsymbol{C},\boldsymbol{D}^{\prime}$) are also sufficient time-dependent adjustment set according to Definition 2 and
    $\sigma^{2}_{\boldsymbol{a},\boldsymbol{B},\boldsymbol{C}}(P)\;\;\geq\;\;\sigma^{2}_{\boldsymbol{a},\boldsymbol{G}^{\prime},\boldsymbol{B},\boldsymbol{C},\boldsymbol{D}^{\prime}}(P)\;\;\geq\;\;\sigma^{2}_{\boldsymbol{a},\boldsymbol{G},\boldsymbol{B},\boldsymbol{C},\boldsymbol{D}}(P)\;\;\geq\;\;\sigma^{2}_{\boldsymbol{a},\boldsymbol{G},\boldsymbol{B}^{\prime},\boldsymbol{C},\boldsymbol{D}^{\prime}}(P)\;\;\geq\;\;\sigma^{2}_{\boldsymbol{a},\boldsymbol{G},\boldsymbol{C}}(P),\ \ and$
    $\sigma^{2}_{\Delta,\boldsymbol{B},\boldsymbol{C}}(P)\;\;\geq\;\;\sigma^{2}_{\Delta,\boldsymbol{G}^{\prime},\boldsymbol{B},\boldsymbol{C},\boldsymbol{D}^{\prime}}(P)\;\;\geq\;\;\sigma^{2}_{\Delta,\boldsymbol{G},\boldsymbol{B},\boldsymbol{C},\boldsymbol{D}}(P)\;\;\geq\;\;\sigma^{2}_{\Delta,\boldsymbol{G},\boldsymbol{B}^{\prime},\boldsymbol{C},\boldsymbol{D}^{\prime}}(P)\;\;\geq\;\;\sigma^{2}_{\Delta,\boldsymbol{G},\boldsymbol{C}}(P).$
4.1 Optimal sufficient time-dependent adjustment set
We now extend a reasoning similar to Rotnitzky and Smucler[5] and Henckel et al [4] to the longitudinal case to demonstrate how to obtain an optimal sufficient time-dependent adjustment set. Our process is based on the subgraphs $\mathcal{G}_{t}$ built as described in Algorithm 1. For each subgraph $\mathcal{G}_{t},t=T,\dots,0$, we define the causal nodes $\mbox{cn}(A_{t},Q_{t},\mathcal{G}_{t})$ as the set of all nodes in $\mathcal{G}_{t}$ (excluding $A_{t}$) that lie on a causal path from $A_{t}$ to $Q_{t}$ with $Q_{T}=Y$ and $\mathcal{G}_{T}=\mathcal{G}$ the initial DAG. The corresponding forbidden set is given by $\mbox{forb}(A_{t},Q_{t},\mathcal{G}_{t})=\mbox{de}(\mbox{cn}(A_{t},Q_{t},\mathcal{G}_{t}))\cup\{A_{t}\}$; the forbidden set thus includes $A_{t}$ as well as the descendants of the causal nodes. An optimal adjustment set for time $t=0,\dots,T$, denoted $\boldsymbol{O}_{t}(A_{t},Q_{t},\mathcal{G}_{t})$, or simply as $\boldsymbol{O}_{t}$ whenever no confusion can arise, is then defined as:
[…]
An optimal sufficient time-dependent adjustment set is obtained by combining these sets: $\boldsymbol{\bar{O}}=(\boldsymbol{O}_{0},...,\boldsymbol{O}_{T})$.
The set $\boldsymbol{O}_{t}$ is an optimal adjustment set at time $t$ in the sense that it is a sufficient adjustment set for estimating the effect of $A_{t}$ on $Q_{t}$ and yielding estimators with the lowest possible asymptotic variance among all sufficient adjustment sets. This is established in Lemmas E.4 and E.5 of Henckel et al [4], as well as Corollary 8 in Rotnitzky and Smucler [5]. Because, as we argued in Section 3.2, the problem of estimating a time-dependent counterfactual mean can be seen as a sequence of estimation of point-treatment counterfactual means, it follows that $\boldsymbol{\bar{O}}$ is also a sufficient time-dependent adjustment set according to Definition 2. By composition, $\boldsymbol{\bar{O}}$ further yields estimators with the smallest possible variance among all time-dependent adjustment sets. Indeed, for any $\boldsymbol{Z}$ that is a sufficient time-dependent adjustment set according to Definition 2, we can use Theorem 4.3 to show that the asymptotic variance of estimators based on $\bar{\boldsymbol{O}}$ is smaller or equal than the variance of estimators based on $\boldsymbol{Z}$. To show this, for all t=0,,Tt=0,...,T, let $\boldsymbol{C}_{t}=\boldsymbol{O}_{t}\cap\boldsymbol{Z}_{t}$, $\boldsymbol{G}_{t}=\boldsymbol{O}_{t}\backslash\boldsymbol{C}_{t}$, $\boldsymbol{B}^{\prime}_{t}=\boldsymbol{Z}_{t}\backslash\boldsymbol{C}_{t}$, $\boldsymbol{D}_{t}=\emptyset$, and choose any $\boldsymbol{B}_{t}\supseteq\boldsymbol{B}^{\prime}_{t}$ disjoint of $(\boldsymbol{G}_{t}$, $\boldsymbol{C}_{t})$ and such that $(\boldsymbol{G},\boldsymbol{B},\boldsymbol{C})$ is a sufficient adjustment set according to Definition 1. It can then be verified that conditions (7), (8) (9) and (10) are met, which implies the desired result. Consequently, we have the following Theorem:
Theorem 4.4.
Consider a DAG $\mathcal{G}$ with vertex set $\boldsymbol{V}$. Let $\boldsymbol{A}\subset\boldsymbol{V}$ and $Y\in\boldsymbol{V}\setminus\boldsymbol{A}$. For $t=T,\dots,0$, define $\mathcal{G}_{t}$ as the subgraph of $\mathcal{G}$ constructed by applying Algorithm 1. Then
$\boldsymbol{\bar{O}}=(\boldsymbol{O}_{0},...,\boldsymbol{O}_{T})$
is a sufficient time-dependent adjustment set according to Definition 2 and, for all $P$ and all sets $\boldsymbol{Z}$ that are sufficient time-dependent adjustment sets according to Definition 2,
$\sigma^{2}_{\boldsymbol{a},\boldsymbol{Z}}(P)-\sigma^{2}_{\boldsymbol{a},\boldsymbol{\bar{O}}}(P)\geq 0.$
$\boldsymbol{\bar{O}}$ is thus an optimal sufficient time-dependent adjustment set.
4.2 Numerical illustrations
@@
We now provide two numerical illustrations of our theoretical results. We first return one last time to Example 1 to illustrate how our theorems can be used in practice to identify a time-dependent adjustment set yielding an estimator with reduced variance or an optimal adjustment set. Next, we revisit Example 4 of Rotnitzky and Smucler (2020) [5]. In this example, an optimal adjustment set cannot be identified based on the causal graph alone when considering Definition 1 and associated lemmas and theorem. Opposingly, our extended Definition 2 and associated theorems resolve this problem.
@@
Table 1 lists all time-dependent adjustment sets according to Definition 2. Note that only adjustment sets 1-9 satisfy Definition 1. Based on Theorem 4.3, the asymptotic variance of estimators based on sets 5 and 14 can be compared, for example. Let $\boldsymbol{K}=(\boldsymbol{K_{0}}=X_{02},\boldsymbol{K_{1}}=X_{12},A_{0},X_{02})$. Because
$A_{1}$ $\perp\!\!\!\perp_{\mathcal{G}}\emptyset\mid X_{02},X_{12},A_{0}$
$A_{0}$ $\perp\!\!\!\perp_{\mathcal{G}_{0}}\emptyset\mid X_{02}$
$Y$ $\perp\!\!\!\perp_{\mathcal{G}}X_{02}\mid X_{12},A_{0},A_{1}$
$X_{02}$ $\perp\!\!\!\perp_{\mathcal{G}_{0}}\emptyset\mid X_{02},A_{0}$
[…]
$X_{01}$ $\sim TN(\mu=0,\sigma^{2}=1,\min=-2,\max=2)$
$X_{02}$ $\sim X_{01}+TN(\mu=0,\sigma^{2}=1,\min=-2,\max=2)$
$A_{0}$ $\sim Bernouilli(p=\text{expit}(X_{01}))$
$X_{11}$ $\sim A_{0}+X_{02}+TN(\mu=0,\sigma^{2}=1,\min=-2,\max=2)$
$X_{12}$ $\sim X_{11}+TN(\mu=0,\sigma^{2}=1,\min=-2,\max=2)$
$A_{1}$ $\sim Bernouilli(p=\text{expit}(A_{0}+X_{02}+X_{11}))$
$Y$ $\sim A_{1}+A_{0}+X_{12}+TN(\mu=0,\sigma^{2}=1,\min=-2,\max=2),$
where $TN$ refers to a truncated normal distribution (in order to avoid practical positivity violations) and $\text{expit}(\cdot)=\exp(\cdot)/(1+\exp(\cdot))$ is the inverse of the logit function. The number of replications was chosen to ensure that the Monte Carlo standard error for the standard deviation was $\approx$ 0.001, which is sufficiently small to accurately identify the adjustment set with the lowest variance. We estimated $\mathbb{E}(Y^{11})$ using a pooled longitudinal targeted maximum likelihood estimator, which is based on the nonparametric efficient influence function $\psi_{P,\boldsymbol{a}}(\boldsymbol{Z};\mathcal{G})$. This was achieved in R v4.3.2 using the ltmle package v2.0.0 and generalized linear models to model both the outcome and exposure.
@@
Adjustment set $\boldsymbol{Z_{0}}$ $\boldsymbol{Z_{1}}$ Monte Carlo
@@
As a second example, we revisit Example 4 from [5]. The associated causal graph is represented in Figure 3. We assume the following time (or topological) ordering of variables $\{H,A_{0},R,Q,A_{1},Y\}$. In this example, the optimal sufficient adjustment set cannot be uniquely identified from the causal graph alone when considering Definition 1. Indeed, as illustrated in [5], the optimal sufficient adjustment set is either 1 or 8 depending on the data-generating equations, more specifically on the relative strength of the $H\rightarrow R\rightarrow Q$ and $H\rightarrow A_{1}$ pathways.
@@
Table 2 lists all sufficient adjustment sets according to Definition 2. Only sets 1-11 satisfy Definition 1. Using Theorem 4.3, it can be found that estimators based on set 24 have a lower variance than those based on either sets 1 or 8. We first compare sets 8 and 24 by applying Theorem 4.3, letting $\boldsymbol{K}=(\boldsymbol{K_{0}}=H,\boldsymbol{K_{1}}=\{A_{0},Q,H\})$, noticing that
$A_{1}$ $\perp\!\!\!\perp_{\mathcal{G}}\emptyset\mid A_{0},H,Q$
$A_{0}$ $\perp\!\!\!\perp_{\mathcal{G}_{0}}H$
$Y$ $\perp\!\!\!\perp_{\mathcal{G}}A_{0},H\mid Q,A_{1}$
$Q_{0}$ $\perp\!\!\!\perp_{\mathcal{G}_{0}}\emptyset\mid H,A_{0}$
[…]
To illustrate empirically these results, we generated 20,000 datasets of size $n$ = 1,000 according to the following equations
@@
Adjustment set $\boldsymbol{Z_{0}}$ $\boldsymbol{Z_{1}}$ Monte Carlo SD
@@
1 $\emptyset$ $\{A_{0},Q\}$ 0.146 0.199
2 $\emptyset$ $\{A_{0},R\}$ 0.175 0.222
3 $\emptyset$ $\{A_{0},H\}$ 0.214 0.248
4 $\emptyset$ $\{A_{0},Q,R\}$ 0.154 0.203
5 $\emptyset$ $\{A_{0},Q,H\}$ 0.165 0.207
6 $\emptyset$ $\{A_{0},R,H\}$ 0.191 0.229
7 $\emptyset$ $\{A_{0},Q,R,H\}$ 0.165 0.207
@@
12 $\emptyset$ $Q$ 0.139 0.194
13 $\emptyset$ $R$ 0.158 0.213
14 $\emptyset$ $\{Q,R\}$ 0.143 0.197
15 $\emptyset$ $\{Q,H\}$ 0.165 0.208
16 $\emptyset$ $\{R,H\}$ 0.191 0.229
17 $\emptyset$ $\{Q,R,H\}$ 0.165 0.208